OpenStudy (dtan5457):

Find the general solutions of 3tanx-sqrt of 3=0

OpenStudy (dtan5457):

@ganeshie8

ganeshie8 (ganeshie8):

solve tanx first

ganeshie8 (ganeshie8):

\[3\tan x -\sqrt{3 }=0\] \[\tan x = \dfrac{\sqrt{3 }}{3}\]

ganeshie8 (ganeshie8):

\[3\tan x -\sqrt{3 }=0\] \[\tan x = \dfrac{\sqrt{3 }}{3} = \dfrac{1}{\sqrt{3}}\] \[x = ?\]

OpenStudy (dtan5457):

is cos=1, sin=sq rt of 3?

OpenStudy (dtan5457):

@ganeshie8

OpenStudy (dtan5457):

oh, tan is 30 degrees.